OpenStudy (anonymous):

Joe climbs a 30 inch ladder to the top of a slide. Find the length of the slide, to the nearest hundredth, if the distance from the base of the ladder to the bottom of the slide is 60 inches. Note: The ladder forms a right angle with the ground

OpenStudy (anonymous):

help????

OpenStudy (mathstudent55):



OpenStudy (anonymous):

what is that supposed to mean? The triangle?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

I still don't get what that is supposed to mean

OpenStudy (anonymous):

where is magbak at??????????????????????????????????????????

OpenStudy (mathstudent55):

The first picture represents the problem. There is a 30-in. ladder that leads to the top of a slide. The bottom of the ladder is 60 in. from the bottom of the slide. The ladder makes a right angle with the ground.

OpenStudy (mathstudent55):

The second picture shows the math you need to solve the problem.

OpenStudy (anonymous):

ok? soo.......

OpenStudy (mathstudent55):

\(30^2 + 60^2 = s^2\) s = length of slide

OpenStudy (anonymous):

ok I'm going to try it.

OpenStudy (mathstudent55):

Great. Square 30, square 60, add the squares together. then take the square root of the sum.

OpenStudy (anonymous):

I got it right!!!!!!!! THANKS!!!!!

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

800 grams is how much kg?

OpenStudy (anonymous):

help?

OpenStudy (mathstudent55):

Do you know how many grams are in a kg?

OpenStudy (drummerboyrevived):

A^ 2+ b^2=c^2